OpenStudy (anonymous):

The specific heat of copper metal is 0.385 J/(g × °C). How much energy must be added to a 35.0 gram sample of copper to change the temperature from 20.0°C to 65.0°C?

OpenStudy (anonymous):

q = mC∆T when q = heat m = mass of substance = 35.0 grams C = 0.385 J/g*C ∆T = change in temperature = 65C - 20C = 45C so q = (35.0 g)*(0.385 J/g*C)*(45C) = 606